Solve the radical equations and check all proposed solutions
√(28 x+28)=x+8

Answer by Cromlix(4381)

√(28 x+28)=x+8
Square both sides
28x + 28 = (x + 8)^2
28x + 28 = x^2 + 16x + 64
Collect like terms
x^2 - 12x + 36 = 0
(x - 6)(x - 6) = 0
x = 6
Hope this helps.
:-)